Commissioning Notes for 20080415 UT

Observers: J. Hill, R. Green, I. Foppiani and a small pod of programmers in Tucson
Telescope Operator: A. Ceranski

Summary

Another good night. Photometric, but worse seeing than last night. Had to reboot the irtc computer to make the socket connections work at the proper speed. Tested some more software. Measured the position angle on a field in M3, but had problems with MCSPU when we tried to adjust it. *Rotated the coordinate system of M2 to match M1.* Lots of stupid problems with IRAF intermittently not being able to create pipes on lbtmu05. Got enough pointing stars to make a decent IRTC pointing model. Got cubes of 15 pupil images averaged in IRAF and into EF, but Richard could not convince EF to analyze them. Almost have a script to take cubes of intra/extrafocal pupil images automatically, but this is held up by an apparent problem with OSS returning to irs before M2 hexapod is finished moving. Had the problem with MCSPU losing polynomials for 1 sec near the end of the night, but it ran OK until 4AM.

Afternoon Activities

HBS seems under control if not fixed. Kevin and Brian think they found that coolant pump #1 is weak. We still had to modify the lookup table because of stratification problems in the tank.

Had a power outage about an hour before sunset, and learned about some new things (Rotator MCR) that can be turned off by a power outage. Ran on generators all night, expect a return outage in the morning.

Wrote a new cookbook for creating a collimation model based on some old emails. Started the cookbook for processing LBT images with EF.

Nighttime Operations

(Times in UT)

07:10 Open dome. Things were a bit slow because of a power outage where we switched to generator involuntarily in late afternoon.

Pointing model is irtc.20080414.ptmod (mask5)

Preset to M67

Problems with irs failing on socket connection to irtc with every GetImage command. Increasing timeout even to 1000 microsec doesn't help. Italo also sees it slow from his laptop.

03:15 Reboot the irtc computer. That cured the problem. timeout is back to 300 microsec. Lost about 30 minutes. After this, socket errors were relatively rare - only a handful during the night.

03:27 Spiral searching on the M67 field, found a nice star (random) near 40,60, and another at -30,90

03:37 Reenable SX actuator 328 to remove fairly gross astigmatism.

03:40 Moved M1 +1 mm, and a coordinated move of M2 -1 mm (reversed) to get M1 off the -Y limit at EL=65. Tonight's focus at EL=63 which is different from the collimation lookup table is at:
M1 SX Global Offsets: X=-0.2, Y=+0.75, Z=0.1, RX=0, RY=0
M2 SX Global Offsets: X=5.5, Y=2.5, Z=0, RX=-75, RY=-280

03:58 Preset to M3 13 42 11.23 +28 22 31.6, start spiral search. Found star at -60,30

ERROR on line 158: cannot open `uparm$pipe24451' for writing. flpr wouldn't cure it. Had to launch a new IRAF session. Had problems with this sporadically through the night. Sometime flpr will cure it, other times we have to start a new IRAF session.

Focus apparently shifted 30 microns in 10 minutes. ?? Seeing is running 1.5 - 2 arcsec, so much worse than last night.

04:23 pointto WT10_277 failed because rotator wasn't running. Aaron gets his first try at turning on the rotator, but it doesn't work. MCR relay on the Rotator Power Box in the Upper Right treehouse had tripped during the power failure. Lost 20 minutes

Started new version of PCS.

04:47 pointto WT10_277, spiral search found at 0,50 from irtc.20080401

*04:58 Pointing image on WT10_277 (but using old pointing model)

0500 Pointing model to irtc.20080414 (mask5), then spiral search again. Found at -60,10

IRAF pipe error again

*05:09 Another pointing image on WT10_277 at EL=66.8 at offset -60,10

Iterated pointing model zeropoints to center the star for offset 0 0 using IA=+164 IE=-30

*05:24 Pointing image on WT10_277 at EL=69.5 at offset=0,2 with new pointing model irtc.20080415.ptmod The first time the pointing model has placed any star on the chip!

05:26 pointto WT10_243, no star, but spiral search found at -25, -27

*05:30 Pointing image on WT10_243 at EL=58.5

Position Angle Measurement

05:32 Preset back to M3, and we see more stars in the field this time. Doug is turning on the AGw from Tucson.

On M3, PA=0 ===> Rotangle 188

takepic 2 offset 0 5 takepic 2 offset 5 0 takepic 2 smeared 05:42 takepic 2 offset 0 0 takepic 2

Resend preset takepic 2

Preset with PA=180, rotator =8 deg takepic 2 smeared takepic 2 smeared recollimate takepic 2

RFG measures that the position angle is off by 45 degrees.

06:02 MCSPU command "roffset lfbg 162000" (rotangle moves from 188 to 232 deg) offset 0 5, but that made N-S move left-right.

06:08 MCSPU command "roffset lfbg -162000" (twice, the offset is apparently relative) But,,,,,,,MCSPU doesn't seem to accept negative values. This is Issue #1265.

06:21 pointto WT10_224 H=6.142, spiral search, found at -25 5

Open left rear vent door

06:27 Pointing Image of WT10_224 at EL=61.9

*06:31 Pointing Image of WT10_224 at EL=61.8 with slightly better collimation

Cubes of Test Pupils for EF analysis

Doug says we should aim for 100 pixel pupil diameters for the EF analysis. The corresponds to extra/intra-focal offsets of -+0.6 mm on either M1 or M2.

20080415031629/ _00007 - sequence of 15 2sec exposures of intrafocal pupil followed by takepic 2

Note that M1 moves nicely in Z for the first 80% of its 1.2 mm move, then deviates by 15 arcsec at the very end. Same behaviour in reverse so there is kind of a position hysteresis loop. Chris thinks this is because of the uncalibrated hardpoint in SX.

20080415031629/ _00008 - sequence of 15 2sec exposures of intrafocal pupil 06:54 followed by takepic 2

20080415031629/ _00009 - sequence of 15 2sec exposures of intrafocal pupil 06:57 followed by takepic 2

20080415031629/ _00010 - sequence of 16 2sec exposures on blank sky

More Pointing Data

07:14 pointto WT10_294

07:14 still another IRAF pipe error: ERROR on line 306: cannot open `uparm$pipe25987' for writing pointto (sourceID=WT10_294)

07:19 and another IRAF pipe error

*07:27 Pointing image of WT10_294 at EL81.6

07:30 pointto WT10_222, but spiral search didn't find star

08:02 pointto WT10_222 again, and try a larger spiral search. No star in +- 2 arcmin.

08:11 pointto WT10_279, found at -18,10 (just outside original field)

*08:20 Pointing Image of WT10_279 at EL=76.3

08:21 pointto WT10_209, found at -52,12

*08:31 Pointing Image of WT10_209 at EL=57.9

Installing new version of OSS that reverses X and Y signs of the hexapod coordinates so they match the directions of M1. We have also reversed the signs of alpha, beta so the operation was equivalent to rotating the M2 coordinate system 180 deg around the optical axis. Nobody is yet willing to make a claim about whether the original signs of alpha and beta were correct.

08:55 pointto WT10_213, found at -90,-60

*09:18 Pointing image of WT10_213 at EL=49.9

Italo took some dark frames here while John was making soup.

Meanwhile at the other end of the room, Richard has succeed in averaging the 15 images in an IRTC image cube with IRAF!

09:30 pointto WT10_411, not found on the first spiral.

Change pointing model to irtc.20080413.ptmod

09:41 pointto WT10_411, found at, -73,-48 but has a faint companion ~8 arcsec away

*09:47 Pointing Image of WT10_411 at EL=37.3

09:48 pointto WT10_345, found at -58,-27

*09:57 Pointing image for WT10_345 at EL=64.9

09:59 pointto WT10_211, found at 4,70

*10:06 Pointing image for WT10_345 at EL=57.5

10:08 pointto WT10_247, found at -8,68

*10:13 Pointing image for WT10_247 at EL=69.7

10:14 pointto WT10_192, found at 40, 68

*10:19 Pointing image for WT10_192 at EL=46.7

10:21 pointto WT10_139, found at 60,75

*10:25 Pointing image for WT10_139 at EL=34.2

10:28 pointto WT10_092, found at 53,97

*10:38 Pointing image for WT10_092 at EL=21.9

10:40 pointto WT10_094, found at 23,95 and a second star at 14,105 (maybe slightly brighter)

10:52 Pointing image for WT10_094a at EL=22.2

10:55 Pointing image for WT10_094b at EL=22.3

10:56 pointto WT10_332, found at -57, -25

*11:07 Pointing image for WT10_332 at EL=67.2

11:08 pointto WT10_369, found at19,-61
04:09 MST El and Az lost polynomials, and again at 04:14, and again at 04:18

*11:30 Pointing image for WT10_369 at EL=72.7

11:35 inex experiment (bad exposure times)

11:40 intra sequence of 15 sky_00022

11:42 extra sequence of 15 sky_00023

11:43 sky sequence of 15 sky_00024

offset 30 -58

intra sequence direct to Repository inex_00025_cube.fit

extra sequence direct to Repository inex_00026_cube.fit

11:50 first test of inex_seq.cl but problems with the focus position

offset 38 -54

11:57 inex_seq.cl

12:00 inex_seq.cl with sky, but OSS got all confused near the end ??

12:03 Stop and restart OSS

offset 42 -50

12:04 inex_seq.cl BUT takepic is starting before moveOPE finishes moving the secondary

This script reports done before the secondary has finished moving.
moveOPE M2 X=0.0 Y=0.0 Z=-0.6 RX=0.0 RY=0.0 RZ=0.0 ABS+ Magic-

This irc command reports done instantly before secondary has finished moving
irc Move 0. 0. -0.6 0. 0. 0. 4 ABS M2 0 left
This is Issue #1266.

irs said: [I] Move command status: SUCCESS
[I] 0
[I] left PSF Move complete

The corresponding irc command to move M1 correctly waits for the mirror to finish moving.

12:12 Close dome

We are leaving telescope with all M2 and M3 swing arms deployed, and all PF swing arms retracted. This is for swing arm commissioning vibration measurements. Aux drives are engaged, so it is ready to drive to horizon even if still parked at zenith.

Two log files in /home/telescope/
-rw-r--r-- 1 telescope lbt 687228 Apr 14 21:22 IRTCpointingmodellog4.14.08
-rw-r--r-- 1 telescope lbt 3352227 Apr 15 06:07 IRTC2pointingmodellog4.14.08

-- JohnHill - 15 Apr 2008
Topic revision: r7 - 19 Jan 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback